OpenStudy (anonymous):

Solve the system using an algebraic approach 3x + 4y + z = 24 5x - 5y - z = 4 5x + y + 4z = 8

OpenStudy (anonymous):

one guy just asked the same problem!

OpenStudy (texaschic101):

3x + 4y + z = 24 5x - 5y - z = 4 ----------------add 8x - y = 28 5x - 5y - z = 4 -->(4) 5x + y + 4z = 8 ------------- 20x - 20y - 4z = 16 (result of multiplying by 4) 5x + y + 4z = 8 ------------add 25x - 19y = 24 8x - y = 28 --->(-19) 25x - 19y = 24 -------------- -152x + 19y = - 532 (result of multiplying by -19) 25x - 19y = 24 -------------add - 127x = - 508 x = -508/-127 x = 4 8x - y = 28 8(4) - y = 28 32 - y = 28 32 - 28 = y 4 = y 5x - 5y - z = 4 5(4) - 5(4) - z = 4 20 - 20 - z = 4 -z = 4 z = -4 check.. 3x + 4y + z = 24 3(4) + 4(4) - 4 = 24 12 + 16 - 4 = 24 28 - 4 = 24 24 = 24 (correct) solution is : x = 4, y = 4, and z = -4

OpenStudy (anonymous):

Thank you!

OpenStudy (texaschic101):

your welcome....elimination is not really that hard, just time-consuming